# SuperEvent: Cross-Modal Learning of Event-based Keypoint Detection for SLAM

### v1

@@ Abstract
Event-based keypoint detection and matching holds significant potential, enabling the integration of event sensors into highly optimized Visual SLAM systems developed for frame cameras over decades of research. Unfortunately, existing approaches struggle with the motion-dependent appearance of keypoints and the complex noise prevalent in event streams, resulting in severely limited feature matching capabilities and poor performance on downstream tasks. To mitigate this problem, we propose SuperEvent, a data-driven approach to predict stable keypoints with expressive descriptors. Due to the absence of event datasets with ground truth keypoint labels, we leverage existing frame-based keypoint detectors on readily available event-aligned and synchronized gray-scale frames for self-supervision: we generate temporally sparse keypoint pseudo-labels considering that events are a product of both scene appearance and camera motion. Combined with our novel, information-rich event representation, we enable SuperEvent to effectively learn robust keypoint detection and description in event streams. Finally, we demonstrate the usefulness of SuperEvent by its integration into a modern sparse keypoint and descriptor-based SLAM framework originally developed for traditional cameras, surpassing the state-of-the-art in event-based SLAM by a wide margin. Source code and multimedia material are available at smartroboticslab.github.io/SuperEvent.
@@ 1 Introduction
Figure 2: Data processing pipeline of SuperEvent: For training, pseudo-labels are generated using a frame-based detector and matcher on the gray-scale frame pairs. We generate spatially synchronized MCTS at the same timestamps and feed them to SuperEvent. The network predictions are compared to the pseudo-labels and the network weights are optimized using backpropagation. During inference, the network predictions can be used to detect keypoints and match their descriptors using the event stream only.
@@
Event cameras offer exciting advantages over their frame-based counterparts, such as an increased temporal resolution, little motion blur, and a high dynamic range. These properties promise to enhance robustness for robot estimation and perception tasks under fast motion and uncontrollable lighting conditions. Processing the sparse and asynchronous output of event cameras, however, requires fundamentally different algorithms than traditional frames. The decades-long research edge in frame-based vision resulted in most hardware and software being highly optimized for frame processing. A common approach for tasks like Structure-from-Motion, Visual Odometry (VO), Simultaneous Localization and Matching (SLAM), or place recognition is to detect and match keypoints across frames to gain a geometric scene understanding [39, 7]. Consequently, a robust keypoint detector and descriptor working on the event stream would unlock the potential of integrating event cameras into extensive and highly-developed algorithms originally developed for frame-based vision.
@@
While several authors already proposed event-based approaches to detect keypoints [47, 9, 10, 31, 32, 18], their matching abilities still show limitations. Most employ nearest-neighbor matching in pixel space, which restricts their usability to visually simple scenes. In an attempt to overcome this issue, more recent approaches learn descriptors in a data-driven fashion. To generate ground truth labels – inspired by frame-based methods [15] – they warp their event representations with a random homography and penalize non-consistent model predictions. This approach forces the models to learn descriptors to match visually similar keypoints. However, it fundamentally ignores the true nature of the event generation process that conflates camera motion with scene appearance and geometry: in essence, matches are generated only from one original set of events per keypoint, thereby producing only limited and potentially also unrealistic training data. We hypothesize that this simplification ultimately results in unstable matching.
@@
In contrast, we propose training our model to predict descriptors for the keypoint correspondences at different points in time – therefore training matches with different and actually recorded rather than warped events. Since generating keypoint and match labels manually is intractable, we design a scalable method to exploit the great capabilities of frame-based vision models: generating pseudo-labels in aligned frames and then using them to train our model’s predictions. An example is shown in Figure 1.
@@
Contributions: (1) We propose a scalable method to generate training data for interest point detection and matching, leveraging robust existing frame-based keypoint detectors. The resulting data incorporates the real temporal dependence of event generation using real event camera recordings. This enables data-driven models to effectively learn robust keypoint detection and description. (2) We develop the model for event-based vision through key design choices: Our novel MCTS event representation increases the performance of data-driven models by reducing their dependence on motion speed of the scene. A transformer backbone shows to further improve the model’s feature extraction. In our conducted experiments, our approach outperforms SOTA detection and matching by a large margin. We publicly release our code to support future event-based vision research. (3) Showcasing the practicality of our method, we integrate the trained model into an existing stereo VI-SLAM approach developed for frame-based vision. The resulting event inertial SLAM (EI-SLAM) system outperforms SOTA approaches on the commonly used TUM-VIE [34] dataset by a large margin.
@@ 2 Related Work
Frame-based Keypoint Detection: there exists a vast variety of approaches to detecting and describing keypoints in frames. Traditional, hand-crafted methods [27, 61, 45, 5, 6, 59, 40] are still widely used, since they are well studied and usually work robustly with little overhead. More recent, data-driven methods outperform the traditional keypoint detectors with impressive results. SuperPoint [15] is the first method trained with homographic adaption, a self-supervised training technique of warping input images and their detected labels by randomized homographies. SuperPoint’s architecture consists of a shared VGG [62]-backbone with a detector and descriptor head, predicting descriptors in a discrete grid which is interpolated to full resolution. The authors later publish the extension SuperGlue [60], a graph neural network that improves the descriptor matching. Following works propose different architectures trained with homographic adaption [63, 65, 57, 21].
Event-based Keypoint Detection: handcrafted approaches can be broadly divided into filters and frame reconstruction. Filtering methods work directly on the event stream attempting to find events originating from corners [11, 1]. While their asynchronous nature offers little processing overhead, downstream tasks are required to process asynchronous data. Existing well-developed downstream solutions for frame-based vision therefore require major modifications.
@@
Additionally, these hand-crafted methods usually employ simple nearest-neighbor matching in pixel space to track keypoints over time. This approach is only reliable for simple scenes since cluttered geometry, high variance in depth, and fast motions cause keypoints to become occluded, tightly clustered, or jumping. Also, the developed heuristics struggle with the ubiquitous and complex noise inherent to event cameras. Usually, they require manual parameter tuning for different camera models, datasets, or sometimes even scenes.
@@
For these reasons, EventPoint [31] and SD2Event [18] employ more complex neural networks that additionally predict descriptors. While EventPoint uses the same architecture as SuperPoint [15] and finetunes its pre-trained weights, SD2Event employs agent-based attention to learn detection and description. Both approaches are trained with homographic adaption and varying time window sizes. While this results in rotation-, distortion- and scale-invariant models, the main challenge of keypoint description remains largely unsolved: Because most events result from scene or camera motion, corresponding keypoints vary in visual appearance. Since robust keypoint matching requires similar descriptor pairs, this motion dependence must be considered. However, homographic adaption creates static frame pairs with little to no variance in scene motion within the samples. This contradicts the dynamic and motion-dependent nature of events, resulting in unstable keypoint matching between different timestamps.
@@
A major drawback of the described tracking approaches is their computational overhead when processing each event individually. Also, most methods only track one keypoint, resulting in an almost linear growth of computational cost with increasing number of tracker instances. E.g., RATE tracks only 20 features in real-time for an event stream of 180x240 pixels resolution – which is hardly sufficient for downstream applications such as SLAM. Additionally, these hand-crafted approaches struggle with environmental and hardware changes, as well as the event cameras’ noise.
@@ 3.1 Training Data Generation
However, annotating keypoints in the asynchronous event data stream manually is very time-consuming and does not scale. Therefore, we propose an approach to generate ground truth data automatically. We take advantage of the pixel-synchronized events and frames of iniVation DAVIS [^1] https://inivation.com/ event cameras. Employing a highly robust frame-based keypoint detector and matcher such as SuperPoint [15] + SuperGlue [60] on the gray-scale frames, we can obtain temporally sparse pseudo-labels. To generate a diverse set of training data with reliable pseudo-labels, we process every frame $i$ of a sequence in the following way:
  • Firstly, we compute the distances in pixel-space of all key points to their respective matches in frame $i+1$. We only consider frames $i$ for training where the distances’ median surpasses a threshold $c_{\text{m}}$. A small median distance indicates that the scene is close to static, resulting in insufficient event data for a useful prediction.
@@ 3.2 Multi-channel Time Surface
  • Time surfaces and the Tencode representation depend on the time window length $\Delta t$. Increasing this parameter not only results in more events being considered but also decreases the contrast between the most recent events and past events. For high optical flow, this negatively impacts the detection of edges. Conversely, a smaller time window increases this contrast but includes fewer events which can fail to sufficiently capture edges in areas with little optical flow. Choosing a suitable $\Delta t$ is challenging, as it depends on the camera and scene motion which might be unknown a priori. The bottom part of Figure 3 shows the effect of varying $\Delta t$: Only the time surfaces in the middle clearly display the net on the top right.
@@ 3.3 Network Architecture & Loss
Unlike other models trained with homographic adaption, our sparse pseudo-labels do not allow for dense descriptor training for every pixel or grid cell. As in [15], we only consider a single descriptor (256 channels) per grid cell, ${\mathbf{y}}_{\text{d}_{h_{0},w_{0}}}$ and ${\mathbf{y}}_{\text{d}_{h_{1},w_{1}}}$ (normalized to length one across the channels) with $(h_{0},w_{0})$ and $(h_{1},w_{1})$ denoting the grid indices of tensor 0 and 1. Successful keypoint matching requires similar descriptors for correspondences and distinct descriptors for unrelated keypoints. We measure descriptor similarity with the dot product $d^{h_{0},w_{0}}_{h_{1},w_{1}}={\mathbf{y}}_{\text{d}_{h_{0},w_{0}}}\cdot{\mathbf{y}}_{\text{d}_{h_{1},w_{1}}}$. Our pseudo label $\hat{y}^{h_{0},w_{0}}_{h_{1},w_{1}}=1$ indicates that two descriptors at the respective grid cell locations correspond, otherwise $\hat{y}^{h_{0},w_{0}}_{h_{1},w_{1}}=0$. Since the majority of the descriptors do not correspond, the partial loss of descriptor matches is weighted with the constant $c_{\text{d}}$. Mathematically, we formulate the descriptor loss as
@@
Figure 4: SuperEvent network architecture and tensor dimensions: a shared transformer backbone is combined with a detector and a descriptor head. On the bottom left, the components of the Convolution-Attention blocks (Conv. + Attn.) and the VGG blocks are displayed. Activations are omitted for simplicity.
@@ 4 Experiments
In our second experiment, we integrate SuperEvent into the SLAM framework OKVIS2 [39] originally developed for frame-based cameras. We achieve convincing results on the widely used event-SLAM benchmark [34] without extensive tuning, showcasing SuperEvent’s versatility.
For ablation studies verifying the effectiveness of temporal matching and the MCTS event representation, examples of the training data and SuperEvents detections, a detailed description of the pose estimation benchmark, as well as plots of the estimated SLAM trajectories, please refer to our Supplementary Material.
@@ 4.1 Implementation Details
[…]
@@ 4.2 Pose Estimation
A common benchmark for evaluating frame-based keypoint detectors is pose estimation [60, 65, 63, 21]. These works use datasets such as ScanNet [12] and MegaDepth [41] containing various images of the same scene with the associated ground truth camera poses. To create a similar event-based benchmark, we employ sequences with ground truth poses from real-world dynamic datasets. In these sequences, we sample timestamps with uniformly distributed pose changes between 1° and 45°. We choose the upper bound to increase the probability of visual overlap. However, since we do not know the ground truth geometry of the scene, it cannot be guaranteed that every sample contains sufficient information to recover the ground truth pose. Next, we run different approaches to detect and match keypoints in the event stream at the identified timestamps. After outlier removal with Random Sample Consensus (RANSAC) [16], we estimate the camera pose and calculate the error with respect to the ground truth. Finally, we report the area-under-curve (AUC) metric for different thresholds.
@@
The experiment is conducted on two diverse datasets: Firstly, the Event Camera Dataset (ECD) [50] is widely used to evaluate event-based approaches, as it was one of the first publicly available datasets. It is challenging due to the low resolution (180x240) and the noisy output of the iniVation DAVIS240C event camera as well as fast motion changes. Secondly, we use the more recent dataset Event-aided Direct Sparse Odometry (EDS) [28]. In contrast to our training data, this dataset is recorded with a Prophesee Gen 3.1222https://www.prophesee.ai/ event camera with a higher resolution of 480x640 pixels and an over-proportionally increased event rate. Fast, inconsistent camera motions as well as a high dynamic range make pose estimation challenging.
Table 1 shows that SuperEvent outperforms all baselines by a large margin. On the ECD dataset, its advantage is the greatest for the high-precision estimations below 5° error, confirming the effectiveness of our method. In contrast, we found that LLAK detects many keypoints but is unable to track the vast majority of them, and EventPoint’s matching capability is severely limited, possibly due to their static assumptions during training inherently contradicting the dynamic nature of event cameras. Finally, RATE achieves decent results for estimations with up to 20° error, but since it cannot recover from tracking loss, it depends on consistent and precise tracking of most keypoints. As seen in the results, this requirement cannot be stably fulfilled.
@@
Table 2 indicates a slight decrease in SuperEvent’s pose estimation performance on the EDS dataset compared to ECD. As the results of the baseline approaches also decrease, we attribute this effect to the dataset being more challenging. Interestingly, the performance of SuperEvent for 20° precision is still in the same region as for the ECD dataset. Considering the different properties of the Prophesee camera compared to the DAVIS camera employed in the training sequences, these results confirm the generalization capability of SuperEvent.
4.3 Runtime
[…]

### v2

@@ Abstract
Event-based keypoint detection and matching holds significant potential, enabling the integration of event sensors into highly optimized Visual SLAM systems developed for frame cameras over decades of research. Unfortunately, existing approaches struggle with the motion-dependent appearance of keypoints and the complex noise prevalent in event streams, resulting in severely limited feature matching capabilities and poor performance on downstream tasks. To mitigate this problem, we propose SuperEvent, a data-driven approach to predict stable keypoints with expressive descriptors. Due to the absence of event datasets with ground truth keypoint labels, we leverage existing frame-based keypoint detectors on readily available event-aligned and synchronized gray-scale frames for self-supervision: we generate temporally sparse keypoint pseudo-labels considering that events are a product of both scene appearance and camera motion. Combined with our novel, information-rich event representation, we enable SuperEvent to effectively learn robust keypoint detection and description in event streams. Finally, we demonstrate the usefulness of SuperEvent by its integration into a modern sparse keypoint and descriptor-based SLAM framework originally developed for traditional cameras, surpassing the state-of-the-art in event-based SLAM by a wide margin. Source code is available at ethz-mrl.github.io/SuperEvent.
@@ 1 Introduction
Figure 2: Data processing pipeline of SuperEvent: For training, pseudo-labels are generated using a frame-based detector and matcher on the gray-scale frame pairs. We generate spatially synchronized MCTS at the same timestamps and feed them to SuperEvent. The network predictions are compared to the pseudo-labels and the network weights are optimized using backpropagation. During inference, the network predictions can be used to detect keypoints and match their descriptors on the event stream only.
@@
Event cameras offer exciting advantages over their frame-based counterparts, such as an increased temporal resolution, little motion blur, and a high dynamic range. These properties promise to enhance robustness for robot estimation and perception tasks under fast motion and uncontrollable lighting conditions. However, processing the sparse and asynchronous output of event cameras requires fundamentally different algorithms than traditional frames. The decades-long research edge in frame-based vision resulted in most hardware and software being highly optimized for frame processing. A common approach for tasks like Structure-from-Motion, Visual Odometry (VO), Simultaneous Localization and Mapping (SLAM), or Place Recognition is to detect and match keypoints across frames to gain a geometric scene understanding [40, 7]. Consequently, a robust keypoint detector and descriptor working on the event stream would unlock the potential of integrating event cameras into extensive and highly-developed algorithms developed initially for frame-based vision.
@@
While several authors already proposed event-based approaches to detect keypoints [48, 9, 10, 32, 33, 19], their matching abilities still show limitations. Most employ nearest-neighbor matching in pixel space, which restricts their usability to visually simple scenes. In an attempt to overcome this issue, more recent approaches learn descriptors in a data-driven fashion. To generate ground truth labels – inspired by frame-based methods [15] – they warp their event representations with a random homography and penalize non-consistent model predictions. This approach forces the models to learn descriptors to match visually similar keypoints. However, it fundamentally ignores the true nature of the event generation process that conflates camera motion with scene appearance and geometry. In essence, matches are generated only from one original set of events per keypoint, thereby producing only limited and potentially also unrealistic training data. We hypothesize that this simplification ultimately results in unstable matching.
@@
In contrast, we propose training our model to predict descriptors for the keypoint correspondences at different points in time – therefore training with different and actually recorded rather than warped events. Since generating keypoint and match labels manually is intractable, we design a scalable method to exploit the great capabilities of frame-based vision models: generating pseudo-labels in aligned frames and then using them to train our model’s predictions. An example is shown in Figure 1.
@@
Contributions: (1) We propose a scalable method to generate training data for interest point detection and matching, leveraging robust existing frame-based keypoint detectors. The resulting data incorporates the real temporal dependence of event generation using real event camera recordings. This enables data-driven models to effectively learn robust keypoint detection and description. (2) We develop the model for event-based vision through key design choices: Our novel MCTS event representation increases the performance of data-driven models by reducing their dependence on motion speed of the scene. A transformer backbone shows to further improve the model’s feature extraction. In our conducted experiments, our approach outperforms SOTA detection and matching by a large margin. We publicly release our code to support future event-based vision research. (3) Showcasing the practicality of our method, we integrate the trained model into an existing stereo VI-SLAM approach developed for frame-based vision. The resulting event inertial SLAM system outperforms SOTA approaches on the commonly used TUM-VIE [35] and VECtor [18] datasets by a large margin.
@@ 2 Related Work
Frame-based keypoint detection: there exists a vast variety of approaches to detecting and describing keypoints in frames. Traditional, handcrafted methods [28, 61, 46, 5, 6, 59, 41] are still widely used, since they are well studied and usually work robustly with little overhead. More recent, data-driven methods outperform the traditional keypoint detectors with impressive results. SuperPoint [15] is the first method trained with homographic adaptation, a self-supervised training technique of warping input images and their detected labels by randomized homographies. SuperPoint’s architecture consists of a shared VGG [62]-backbone with a detector and descriptor head, predicting descriptors in a discrete grid which is interpolated to full resolution. The authors later publish the extension SuperGlue [60], a graph neural network that improves the descriptor matching. Following works propose different architectures trained with homographic adaptation [63, 65, 22].
Event-based keypoint detection: handcrafted approaches can be broadly divided into filters and frame reconstruction. Filtering methods work directly on the event stream, attempting to find events originating from corners [11, 1]. While their asynchronous nature offers little processing overhead, downstream tasks are required to process asynchronous data. Existing well-developed downstream solutions for frame-based vision therefore require major modifications.
@@
Additionally, these handcrafted methods usually employ simple nearest-neighbor matching in pixel space to track keypoints over time. This approach is only reliable for simple scenes since cluttered geometry, high variance in depth, and fast motions cause keypoints to become occluded, tightly clustered, or to jump. Also, the developed heuristics struggle with the ubiquitous and complex noise inherent to event cameras. Usually, they require manual parameter tuning for different camera models, datasets, or sometimes even scenes.
@@
For these reasons, EventPoint [32] and SD2Event [19] employ more complex neural networks that additionally predict descriptors. While EventPoint uses the same architecture as SuperPoint [15] and finetunes its pre-trained weights, SD2Event employs agent-based attention to learn detection and description. Both approaches are trained with homographic adaptation and varying time window sizes. While this results in rotation-, distortion-, and scale-invariant models, the main challenge of keypoint description remains largely unsolved: Because most events result from scene or camera motion, corresponding keypoints vary in visual appearance. Since robust keypoint matching requires similar descriptor pairs, this motion dependence must be considered. However, homographic adaptation creates static frame pairs with little to no variance in scene motion within the samples. This contradicts the dynamic and motion-dependent nature of events, resulting in unstable keypoint matching between different timestamps.
@@
A major drawback of the described tracking approaches is their computational overhead when processing each event individually. Also, most methods only track one keypoint, resulting in an almost linear growth of computational cost with increasing number of tracker instances. E.g., RATE tracks only 20 features in real-time for an event stream of $180\times 40$ pixels resolution – which is hardly sufficient for downstream applications such as SLAM. Additionally, these handcrafted approaches struggle with environmental and hardware changes, as well as the event cameras’ noise.
@@ 3.1 Training Data Generation
However, manually annotating keypoints in the asynchronous event data stream is very time-consuming and does not scale. Therefore, we propose an approach to generate ground truth data automatically. We take advantage of the pixel-synchronized events and frames of iniVation DAVIS [^1] https://inivation.com/ event cameras. Employing a highly robust frame-based keypoint detector and matcher such as SuperPoint [15] + SuperGlue [60] on the gray-scale frames, we can obtain temporally sparse pseudo-labels. To generate a diverse set of training data with reliable pseudo-labels, we process every frame $i$ of a sequence in the following way:
  • Firstly, we compute the distances in pixel-space of all key points to their respective matches in frame $i+1$. We only consider frames $i$ for training where the distances’ median surpasses a threshold $c_{\text{m}}$. A small median distance indicates that the scene is nearly static, resulting in insufficient event data for a useful prediction.
@@ 3.2 Multi-channel Time Surface
  • Time surfaces and the Tencode representation depend on the time window length $\Delta t$. Increasing this parameter not only results in more events being considered but also decreases the contrast between the most recent events and past events. For high optical flow, this negatively impacts the detection of edges. Conversely, a smaller time window increases this contrast but includes fewer events, which can fail to sufficiently capture edges in areas with little optical flow. Choosing a suitable $\Delta t$ is challenging, as it depends on the camera and scene motion which might be unknown a priori. The bottom part of Figure 3 shows the effect of varying $\Delta t$: Only the time surfaces in the middle clearly display the net on the top right.
@@ 3.3 Network Architecture & Loss
Unlike other models trained with homographic adaptation, our sparse pseudo-labels do not allow for dense descriptor training for every pixel or grid cell. As in [15], we only consider a single descriptor (256 channels) per grid cell, ${\mathbf{y}}_{\text{d}_{h_{0},w_{0}}}$ and ${\mathbf{y}}_{\text{d}_{h_{1},w_{1}}}$ (normalized to length one across the channels) with $(h_{0},w_{0})$ and $(h_{1},w_{1})$ denoting the grid indices of tensor 0 and 1. Successful keypoint matching requires similar descriptors for correspondences and distinct descriptors for unrelated keypoints. We measure descriptor similarity with the dot product $d^{h_{0},w_{0}}_{h_{1},w_{1}}={\mathbf{y}}_{\text{d}_{h_{0},w_{0}}}\cdot{\mathbf{y}}_{\text{d}_{h_{1},w_{1}}}$. Our pseudo label $\hat{y}^{h_{0},w_{0}}_{h_{1},w_{1}}=1$ indicates that two descriptors at the respective grid cell locations correspond, otherwise $\hat{y}^{h_{0},w_{0}}_{h_{1},w_{1}}=0$. Since the majority of the descriptors do not correspond, the partial loss of descriptor matches is weighted with the constant $c_{\text{d}}$. Mathematically, we formulate the descriptor loss as
@@
Figure 4: SuperEvent network architecture and tensor dimensions: a shared transformer backbone is combined with a detector and a descriptor head. The components of the Convolution-Attention blocks (Conv. + Attn.) and the VGG blocks are displayed on the bottom left. Activations are omitted for simplicity.
@@ 4 Experiments
Table 1: Inference times of our compiled PyTorch model on GPU for common event-stream resolutions (cropped to compatible shape) using float16 precision (same accuracy as float32 in our experiments). NMS requires an additional $0.4\text{\,}\mathrm{m}\mathrm{s}$ on average.
$160\times 240$ $240\times 320$ $480\times 640$ $720\times 1280$
$1.3\text{\,}\mathrm{m}\mathrm{s}$ $2.6\text{\,}\mathrm{m}\mathrm{s}$ $7.8\text{\,}\mathrm{m}\mathrm{s}$ $23.2\text{\,}\mathrm{m}\mathrm{s}$
@@
In our second experiment, we integrate SuperEvent into the SLAM framework OKVIS2 [40] originally developed for frame-based cameras. We achieve convincing results on the widely used event-SLAM benchmarks [35, 18] without extensive tuning, showcasing SuperEvent’s versatility.
For ablation studies, more qualitative examples, a detailed description of the pose estimation benchmark, and a comparison to frame-based matching for high-speed/ HDR scenarios, please refer to our Supplementary Material.
@@ 4.1 Implementation Details
We employ the network architecture as in Section 3.3 with the loss function hyperparameters $c_{\lambda}=10$, $c_{\text{d}}=0.5$, $c_{\text{p}}=1.0$, and $c_{\text{n}}=0.2$. We encode events as a 10-channel MCTS with 5 time window sizes $\Delta t$ logarithmically distributed from $0.001\text{\,}\mathrm{s}$ to $0.1\text{\,}\mathrm{s}$. The detection threshold is set to $c_{\text{s}}=0.01$ and the NMS neighborhood $c_{\text{b}}=2$ pixels.
We evaluate on a desktop PC with an Intel Core i5-13600 processor, an NVIDIA GeForce GTX 4070 GPU, and $32\text{\,}\mathrm{G}\mathrm{B}$ RAM. Inference times are shown in Table 1.
@@ 4.2 Pose Estimation
A common benchmark for evaluating frame-based keypoint detectors is pose estimation [60, 65, 63, 22]. These works use datasets such as ScanNet [12] and MegaDepth [42] containing various images of the same scene with the associated ground truth camera poses. To create a similar event-based benchmark, we employ sequences with ground truth poses from real-world dynamic datasets. We search for uniformly distributed pose changes between 1° and 45°. The upper bound is chosen to increase the probability of visual overlap. However, since we do not know the ground truth geometry of the scene, it cannot be guaranteed that every sample contains sufficient information to recover the ground truth pose. Next, we process the event stream with various approaches to detect and match keypoints. After outlier removal with Random Sample Consensus (RANSAC) [16], we estimate the camera pose and calculate the error with respect to the ground truth. Finally, we report the area-under-curve (AUC) metric for different thresholds.
@@
The experiment is conducted on two diverse datasets: Firstly, the Event Camera Dataset (ECD) [51] is widely used to evaluate event-based approaches, as it was one of the first publicly available datasets. It is challenging due to the low resolution ($180\times 240$), the noisy output of the iniVation DAVIS240C event camera, and fast motion changes. Secondly, we use the more recent dataset Event-aided Direct Sparse Odometry (EDS) [29]. In contrast to our training data, this dataset is recorded with a Prophesee Gen 3.1222https://www.prophesee.ai/ event camera with a higher resolution of $480\times 640$ pixels and an over-proportionally increased event rate. Fast, inconsistent camera motions and a high dynamic range make pose estimation challenging.
Table 2 shows that SuperEvent outperforms all baselines by a large margin. On the ECD dataset, its advantage is the greatest for the high-precision estimations below 5° error, confirming the effectiveness of our method. In contrast, we found that LLAK detects many keypoints but cannot track the vast majority of them. EventPoint’s matching capability is severely limited, possibly due to their static assumptions during training inherently contradicting the dynamic nature of event cameras. Finally, RATE achieves decent results for estimations with up to 20° error, but since it cannot recover from tracking loss, it depends on consistent and precise tracking of most keypoints. As seen in the results, this requirement cannot be stably fulfilled.
@@
Table 3 indicates a slight decrease in SuperEvent’s pose estimation performance on the EDS dataset compared to ECD. Since all approaches perform worse, we attribute this effect to the dataset being more challenging. Interestingly, the performance of SuperEvent for 20° precision is still in the same region as for the ECD dataset. Considering the different properties of the Prophesee camera compared to the DAVIS camera employed in the training sequences, these results confirm the generalization capability of SuperEvent.